# Sphere Packing on a Quantum Computer for Chromatography Modeling

### v1

@@ V Conclusion
In this work, we translated the biopharmaceutical problem of protein filtration to the mathematical problem of sphere packing. We identified three complexity levels of sphere packing. The first level was solved on real quantum hardware, the second was simulated on a classical computer (with a road map to quantum simulations), and the third was shown to be an extension of the first two. The fact that classical algorithms scale exponentially for this NP-hard problem opens up the possibility of quantum advantage in the near to long term. Robust resource scaling of the quantum resources required to solve the problem on a quantum computer was conducted to further support this claim. In the future, we would like to attempt to solve the “heterogeneous circle packing” complexity level on a quantum computer in the near-term and the “heterogeneous sphere packing” complexity in the long term. Altogether, this work provided a base and road map to potential quantum advantage for the problem of protein filtration modeled as sphere packing, which could have profound positive implications for the future of the biopharmaceutical industry and the health and wellness of people everywhere.
@@ VI Acknowledgments
This work was funded by CSL Behring. The authors used Infleqtion’s compilation software, Superstaq, to submit circuits to IQM’s quantum computer, Garnet, through AWS’s Braket. From Infleqtion, we would like to thank Pranav Gokhale for guidance on quantum hardware selection, Caitlin Carnahan for serving as program manager for this project, and Peter Noell for managing coordination between Infleqtion and CSL. Additionally, we acknowledge Cameroon Booker for theoretical insights including the graph formulation for the hardware experiment. Finally, we acknowledge Bao Bach, Rajas Chari, Kabir Dubey, Rohan Mehta, Shivam Mundhra, and Carlo Siebenschuh for proposing the formulation of chromatography as sphere packing in their submission (mentored by Infleqtion’s Victory Omole) at the quantum hackathon Big Q Hack, upon which this work was inspired.

### v2

@@ V Conclusion
In this work, we model a packed chromatography column using a mathematical formulation of sphere packing. We identified three complexity levels of sphere packing. The first level was solved on real quantum hardware, the second was simulated on a classical computer (with a road map to quantum simulations), and the third was shown to be an extension of the first two. Accompanying the quantum hardware results: hyperparameter optimization was conducted to determine the optimal penalty term $\lambda$ for the problem, optimal compilation was implemented to reduce noise, and noisy simulation was conducted to preemptively suggest a number of layers $p$ with which to run QAOA on quantum hardware. Finally, evidence of parameter concentration was found for the problem, suggesting that large, untrainable instances of the problem may still be solvable by QAOA via execution on parameters trained on smaller, trainable, problem instances. The fact that classical algorithms scale exponentially for this NP-hard problem opens up the possibility of quantum advantage in the near to long term. Robust resource scaling of the quantum resources required to solve the problem on a quantum computer was conducted to further support this claim. In the future, we would like to attempt to solve the “heterogeneous circle packing” complexity level on a quantum computer in the near-term and the “heterogeneous sphere packing” complexity level in the long term. Altogether, this work provided a base and road map to potential quantum advantage for the problem of column chromatography modeled as sphere packing, which could have profound positive implications for the future of the biopharmaceutical industry and the health and wellness of the public.
@@ VI Acknowledgments
This work was funded by CSL Behring. The authors used Infleqtion’s compilation software, Superstaq, to submit circuits to IQM’s quantum computer, Garnet, through AWS’s Braket. From Infleqtion, we would like to thank Pranav Gokhale for guidance on quantum hardware selection, Caitlin Carnahan for serving as program manager for this project, and Peter Noell for managing coordination between Infleqtion and CSL. Additionally, we acknowledge Cameron Booker for theoretical insights including the graph formulation for the hardware experiment. Additionally, we acknowledge Bao Bach, Rajas Chari, Kabir Dubey, Rohan Mehta, Shivam Mundhra, and Carlo Siebenschuh for proposing the formulation of chromatography as sphere packing in their submission (mentored by Infleqtion’s Victory Omole) at the quantum hackathon Big Q Hack, upon which this work was inspired. Finally, we acknowledge the work of every scientist that has come before us as “if (we) have seen further, it is by standing on the shoulders of giants” - Sir Isaac Newton.